OpenStudy (anonymous):

Verify the identity. Justify each step. tan (theta) + cot (theta) = 1 / sin (theta) cos (theta)

OpenStudy (anonymous):

tan(theta) = sin(theta) / cos(theta) and cot(theta) = cos(theta) / sin(theta) Substitute these in the above equation and you will prove it right...Any doubt?

OpenStudy (anonymous):

How would I substitute the 'sin'.?

OpenStudy (anonymous):

tanx + cotx = sinx/cosx + cosx/sinx = \[( \sin x ^{2} + \cos ^{2} ) / (\sin x)(\cos x) = 1/(\sin x)(\cos x)\]

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

we know, \[(\sin x) ^{2} + (\cos x) ^{2} = 1\]

zepdrix (zepdrix):